OpenStudy (anonymous):

How long would a 1000 mile trip take if I was traveling at 50 miles per hour for the first 500 miles and 60 miles an hour for the rest of the trip?

OpenStudy (goldphenoix):

How long did it take that person to drive the first 500 miles?

OpenStudy (anonymous):

10?

OpenStudy (goldphenoix):

Good. So keep 10 hours in mind. So you drove 500 miles. Now you need 500 miles left. Since he is driving 60 miles per hour, how long will it take him to drive 500 miles?

OpenStudy (jdoe0001):



OpenStudy (anonymous):

8.3

OpenStudy (jdoe0001):

:)

OpenStudy (goldphenoix):

Good. So you have 8.3 hours and 10 hours. Add that up, and you'll get your answer.

OpenStudy (anonymous):

the anser is 6/5 tho

OpenStudy (anonymous):

i keep on getting 18.3

OpenStudy (jdoe0001):

then you set 8.3 as a rational, 500/60, then simplify

OpenStudy (jdoe0001):

500/50 simplifies already to 10

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

right and \(\bf \cfrac{500}{60} \implies \cfrac{25}{3} \implies 8 \cfrac{1}{3}\)

OpenStudy (jdoe0001):

\(\bf \large 10+\frac{25}{3} \implies \frac{55}{3} \implies 18\frac{1}{3}\)